OpenStudy (anonymous):

Simplify: a. ^33 sqrt x^55 b. ^3 sqrt 64y^105

Nnesha (nnesha):

33 is exponent of what ??

OpenStudy (anonymous):

The formatting of this isn't very clear, are you asking \[a. \sqrt[33]{x^(55)}\] and \[b. \sqrt[3]{64y^105}\]

OpenStudy (anonymous):

\[(\sqrt{x^{55}})^{33}\] is that what you mean

Nnesha (nnesha):

hmm that is 33d root ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

\[\huge\rm \sqrt[33]{x^{55}}\] ??

OpenStudy (anonymous):

\[\sqrt[33]{x^{55}}=x^{\frac{55}{33}}=x^{\frac{5}{3}}=\sqrt[3]{x^5}\]

OpenStudy (anonymous):

\[\sqrt[3]{64y^{105}}=\sqrt[3]{64} \times \sqrt[3]{y^{105}}=4 \times y^{\frac{105}{3}}=4 \times y^{35}=4y^{35}\]

OpenStudy (kohai):

Please close the question if you no longer need help :)